OpenStudy (anonymous):

If a 40 kg gymnast and a 400 kg sumo wrestler each dropped from 1 m above the trampoline, find the final position of each athlete. Assume the trampoline is a simple spring obeying Hooke's law with a k value of 12 000 N/m.

OpenStudy (anonymous):

The force on each of the bodies is measured as F = MA were acceleration is gravity so you can find the force for both the people.

OpenStudy (anonymous):

Using conservation of energy you know that 1 person first begins as potential energy and later it is transformed into Elastic potential energy so\[U_g = U_S\] Where \[U_g = mass \times gravity \times height \] and \[U_s = \frac{ 1}{ 2 } \times K \times X^2\]Where K is the spring constant and X is the distance

OpenStudy (anonymous):

Now to find the final position, set them equal t each other and solve for X